# Convex Formulation of the Zero Emission Vessel Route Planning Problem

### v2

@@ 3.3 Example problem with two variables
$x_{2}\leq 2x_{1}^{-1},\,x_{2}\leq 3.$
3.4 Fitting a log-convex function to data
We consider the problem of modeling data generated by sampling a non-log-convex function or a black-box simulation routine. This situation arises when modeling complex physical phenomena like hull hydrodynamics. Our starting point is the $N$ generated data points
$(x_{i},f_{i})\in\mathbb{R}^{n}_{++}\times\mathbb{R}_{++},\,i=1,\ldots,N.$
If we can fit the data with a monomial or a posynomial, the model encoded by the data becomes available in an analytical log-convex functional form.
The fitting problem concerns fitting a convex function to the data points
$(y_{i},g_{i})\in\mathbb{R}^{n}\times\mathbb{R},\,i=1,\ldots,N,$
which have been obtained by the log-transformation $(y_{i},g_{i})=(\log x_{i},\log f_{i})$.
Convex functions belonging to the class of softmax-linear have been observed to work well in practice with only a small number of terms (Hoburg et al., 2016). This class of functions has the form
$f_{\mathrm{SMA}}(y)=\frac{1}{\alpha}\log\sum_{k=1}^{K}\exp(\alpha(b_{k}+a_{k}^{T}y)),$ (5)
where the parameters are $\alpha>0$, $a_{k}\in\mathbb{R}^{n}$ and $b_{k}\in\mathbb{R}$. The number of user-specified exponential terms is $K$.
The nonlinear least squares problem seeks the parameters $a,b$, and $\alpha$ that minimize the sum of squared residuals
minimize $\sum_{i=1}^{N}(f_{\mathrm{SMA}}(y_{i},a,b,\alpha)-g_{i})^{2}$ (6)
with the decision variables $a,b$ and $\alpha$. Since this is a nonconvex optimization problem, the quality of the fit can depend on the initial guess of the parameters.
The softmax-linear function corresponds to the posynomial constraint
$\sum_{k=1}^{K}\exp(\alpha b_{k})\prod_{i=1}^{n}x_{i}^{\alpha a_{ik}}\leq f^{\alpha},$
where the parameters $a,b$ and $\alpha$ are the minimizers of the least squares problem (6).
4 Transportation network model
All the symbols used in §3 are redefined in this and the following sections.
4.1 Network structure
We adopt the notion of a service, which is a sequence of ports visited in a regular schedule. One vessel or several identical vessels operate a service. A shipping company typically operates several services among a set of ports in a given geographical area (e.g., the Baltic Sea or the Mediterranean Sea). The demands of cargoes between all pairs of ports are known quantities. The vessels can transport different types of cargoes, e.g., cars, trucks, trailers and passengers, as in the roll-on/roll-off passenger ferry segment. The following notation representing the services and cargo demands builds on the liner network design problem presented by Christiansen et al. (2007).
The ports in the network are linearly ordered and numbered from $1$ to $N$. Each service has two end ports $i$ and $j$, where $1\leq i<j\leq N$. A service route starts in $i$ and proceeds outbound to ports with higher and higher numbers until it reaches $j$. At $j$, the route turns around and travels inbound to ports with lower and lower numbers until it reaches $i$. A service with $i$ and $j$ as end ports does not need to call at all ports between $i$ and $j$. However, the port calls are the same on the route’s outbound and inbound legs. Figure 2 illustrates outbound legs of three routes.
4.2 Indexing sets
Figure 2: A route is a sequence of sailing legs between linearly ordered ports from 1 to $N$ and indexed by $i$ and $j$. The figure illustrates a shipping network consisting of three routes constructed from subsets of the same set of $N$ ports. If two or more routes include a pair of ports, multiple services can supply the cargo demand between these ports.
We let $\mathcal{S}$ be the set of services indexed by $s$ and $\mathcal{C}$ the different types of cargoes indexed by $c$. Furthermore, let $\mathcal{N}$ be a totally ordered set of the ports $1$ to $N$ in a linear order, indexed by $i,j,i^{\prime}$ and $j^{\prime}$. The origin port of service $s\in\mathcal{S}$ is $o_{s}$ and the destination port is $d_{s}$. The set $\mathcal{N}_{s}\subseteq\mathcal{N}$ denotes the ports of service $s\in\mathcal{S}$.
$\mathcal{N}^{-}_{is}=\{o_{s},\ldots,i\}\subseteq\mathcal{N}_{s}$
and the ports that follow $i$ is the set
$\mathcal{N}^{+}_{is}=\{i,\ldots,d_{s}\}\subseteq\mathcal{N}_{s}.$
[…]
$\mathcal{L}_{s}=\left\{(i,j):i\in\mathcal{N}_{s}\backslash d_{s},j=\min\mathcal{N}_{is}^{+}\backslash\left\{i\right\}\right\}.$
@@ 4.2 Indexing sets
4.3 Capacity constraints
We let $f_{ijcs}$, for all $(i,j)\in\mathcal{A},c\in\mathcal{C},s\in\mathcal{S}$, denote the decision variable for the quantity of cargo $c$ transported from origin port $i$ to the destination port $j$ by service $s$ during the planning horizon. The transportation capacity of cargo $c$ of a vessel in service $s$ is $f_{cs}^{\text{cap}}$, which is a decision variable governed by constraints specified in the vessel design model introduced in §5. The variable $N^{\text{rt}}_{s}\in\mathbb{Z}_{\geq 1},s\in\mathcal{S}$ denotes the number of whole round-trip voyages that a vessel completes in service $s$ during the planning horizon.
The capacity constraints for the outbound legs are written as standard posynomial constraints
@@ 4.3 Capacity constraints
for all $c\in\mathcal{C},\ s\in\mathcal{S},\,i\in\mathcal{N}_{s}\backslash d_{s}$. For the inbound legs, the capacity constraints take the same form as above, but the departure and arrival ports appear in a reverse order:
@@
4.4 Demand constraints
The demand for cargo $c$ on the arc $(i,j)$ over the planning horizon is given by the fixed parameter $f_{ijc}^{\text{dem}}$ for all $(i,j)\in\mathcal{A},\,c\in\mathcal{C}$. The demand on the arc $(i,j)$ is the upper bound on the total quantity of cargo transported by all the services that include these ports in their routes. We write this demand constraint as
$\sum_{s:(i,j)\in\mathcal{A}_{s}}f_{ijcs}\leq f_{ijc}^{\text{dem}},\quad c\in\mathcal{C},\ (i,j)\in\mathcal{A}.$ (9)
4.5 Availability constraints
The decision variable $v_{ijs}$, for all $(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S}$, denotes the sailing speed from port $i$ to $j$ in service $s$ and the parameter $l_{ij}^{\text{leg}}$ denotes the sailing distance. The sailing time $t_{ijs}^{\text{sea}}$ is the sum of the acceleration time, the deceleration time, and the constant speed sailing time:
$t_{ijs}^{\text{sea}}=\frac{l_{ij}^{\text{leg}}}{v_{ijs}}+\frac{v_{ijs}}{2}\left(\frac{1}{\dot{v}_{s}^{(+)}}+\frac{1}{\dot{v}_{s}^{(-)}}\right),$
where the acceleration and deceleration rates $\dot{v}_{s}^{(-)}$ and $\dot{v}_{s}^{(+)}$ are constant and uniform for each leg in each route.
The cargo loading and unloading times $t_{ijsc}^{\text{cl}}$ and $t_{ijsc}^{\text{cu}}$ at the destination port of the leg $(i,j)$ in service $s$ for cargo $c$ are given as
@@ 4.5 Availability constraints
where $t_{c}^{\text{unit}}$ is the loading and unloading duration of a cargo unit $c$. We let the decision variable $t_{ijs}^{\text{cha}}$ for all $(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S}$ denote the charging time at the destination port of the leg $(i,j)$ of a vessel in service $s$. The docking time $t_{ijs}^{\text{port}}$ is then the time of the port operation that takes the longest:
$t_{ijs}^{\text{port}}=\max\{\max\{t_{ijsc}^{\text{cl}}:c\in\mathcal{C}\},\,\max\{t_{ijsc}^{\text{cu}}:c\in\mathcal{C}\},t_{ijs}^{\text{cha}}\}.$
Using all the above definitions, we write the availability constraints in the following form, accounting for both the outbound and inbound legs:
$\sum_{(i,j)\in\mathcal{L}_{s}}N^{\text{rt}}_{s}(t_{ij}^{\text{sea}}+t_{ji}^{\text{sea}}\,+t_{ij}^{\text{port}}+t_{ji}^{\text{port}})\leq t_{s}^{\text{route}},\,s\in\mathcal{S},$ (10)
where the parameter $t_{s}^{\text{route}}$ is the maximum time a vessel in service $s$ is available during the planning horizon.
4.6 Transportation service level constraint
We require that the transportation network provides a minimum service level:
$U^{\mathrm{network}}\geq U^{\mathrm{min}}.$ (11)
Here we use log-utility, which is a particular concave function with diminishing incremental value of cargo volume shipped in each origin-destination pair and cargo type:
$\begin{split}U^{\mathrm{network}}&=\sum_{(i,j)\in\mathcal{A}}\sum_{s:(i,j)\in\mathcal{A}_{s}}\sum_{c\in\mathcal{C}}\alpha_{ijc}\log f_{ijcs}\\ &=\log\prod_{(i,j)\in\mathcal{A}}\prod_{s:(i,j)\in\mathcal{A}_{s}}\prod_{c\in\mathcal{C}}f_{ijcs}^{\alpha_{ijc}},\end{split}$ (12)
where $\alpha_{ijc}$ is a parameter that weights the cargo flows by their arc lengths and cargo value.
5 Vessel design model
This section describes a coupled system of inequalities that captures the key naval-architectural principles governing vessel design. The inequalities are log-convex compatible, meaning we can manipulate them to the form $1\geq f(x)$, where $f$ is a log-convex function. All the new constraints and the associated design variables hold for each $s\in\mathcal{S}$, but we drop the index for clarity.

### v3

@@ 4.1 Indexing sets
$\mathcal{L}_{s}=\left\{(i,j),(j,i):i\in\mathcal{N}_{s}\backslash d_{s},j=\min\mathcal{N}_{is}^{+}\backslash\left\{i\right\}\right\}.$
@@
4.2 Parameters
The cost of installing one unit of charger power is $C^{\text{cha}}$, and the cost of charging one unit of electrical energy is $C^{\text{el}}$. The onshore power installed in port $i$ is upper bounded to $P^{\text{cha,max}}_{i}$. Port charge is levied in proportion to the ship’s gross tonnage and denoted as $C^{\text{port}}$. The demand and value of a unit of cargo $c$ shipped from port $i$ to $j$ are $q^{\text{dem}}_{ijc}$ and $\alpha_{ijc}$. The constant $U^{\text{min}}$ is the total required network transportation utility. The distance of sailing directly from port $i$ to $j$ is $l^{\text{leg}}_{ij}$. It takes $t^{\text{unit}}_{c}$ units of time to load a unit of cargo $c$. A ship in service $s$ is available at most $t^{\text{route}}_{s}$ time during the planning horizon. Ships sailing on overnight cruises are available for all hours of the day, but in short shuttle services, they are docked overnight.
4.3 Decision variables
We use the following decision variables: the positive integer variable $N^{\text{ship}}_{s}\in\mathbb{Z}_{\geq 1},s\in\mathcal{S}$ gives the number of ships deployed in service $s$. Another integer variable $N^{\text{trip}}_{s}\in\mathcal{F},s\in\mathcal{S}$ gives the total number of roundtrips completed in service $s$ by all the ships (each ship completes $N^{\text{trip}}_{s}/N^{\text{ship}}_{s}$ roundtrips). Variables associated with the sailing leg $(i,j)$ for ship $s$ are the sailing speed $v_{ijs}\in\mathbb{R}_{++},(i,j)\in\mathcal{L}_{s},s\in\mathcal{S}$ and the turnaround time at the destination port $t^{\text{port}}_{ijs}\in\mathbb{R}_{++},(i,j)\in\mathcal{L}_{s},s\in\mathcal{S}$. A ship in service $s$ transports the quantity $q_{ijcs}\in\mathbb{R}_{++},(i,j)\in\mathcal{A}_{s},c\in\mathcal{C},s\in\mathcal{S}$ of cargo $c$ from port $i$ to $j$ during the planning horizon. The only variable associated with shore infrastructure is $P_{i}^{\text{cha}}\in\mathbb{R}_{++},i\in\mathcal{N}$, which is the installed charger power at port $i$. The ship design variables depend on the particular ship design model used. Here we represent them with the $p$-vector $x^{\text{ship}}_{s}\in\mathbb{R}_{++}^{p},s\in\mathcal{S}$.
4.4 Functions
Some quantities are represented as functions of the ship design variables $x^{\text{ship}}_{s}$ and sailing leg speeds $v_{ijs}$. Let $E^{\text{leg}}_{ij}:\mathbb{R}_{++}^{k\times 1}\rightarrow\mathbb{R}_{++}$ denote ship electrical energy consumption sailing from port $i$ to $j$. This leg-dependent formulation accounts for sailing distance. We also define $C^{\text{ship}}:\mathbb{R}_{++}^{k}\rightarrow\mathbb{R}_{++}$ as ship construction and lifetime operation cost excluding energy. Ship capacity of cargo $c$ is $q^{\text{cap}}_{c}:\mathbb{R}_{++}^{k}\rightarrow\mathbb{R}_{++}$ and gross tonnage is $V^{\text{GT}}:\mathbb{R}_{++}^{k}\rightarrow\mathbb{R}_{++}$.
4.5 Model formulation
Table 1: Model notation
Symbol Description Symbol Description
Sets Parameters
$\mathcal{A}$ Origin-destination demand arcs $\alpha_{ijc}$ Utility of a unit of cargo $c$ shipped
$\mathcal{A}_{s}$ Origin-destination demand arcs of service $s$ from port $i$ to $j$
$\mathcal{C}$ Cargoes $C^{\text{cha}}$ Charger installation cost
$\mathcal{F}$ Service frequencies $C^{\text{port}}$ Port charge
$\mathcal{L}_{s}$ Outbound sailing legs in service $s$ $C^{\text{el}}$ Electricity price
$\mathcal{N}$ Ports $d_{s}$ Outbound destination port of service $s$
$\mathcal{N}_{s}$ Ports visited by service $s$ $q^{\text{dem}}_{ijc}$ Demand of cargo $c$ from port $i$ to $j$
$\mathcal{N}_{is}^{–}$ Ports in service $s$ that precede port $i$ $l^{\text{leg}}_{ij}$ Sailing distance from port $i$ to $j$
$\mathcal{N}_{is}^{+}$ Ports in service $s$ that follow port $i$ $o_{s}$ Outbound departure port of service $s$
$\mathcal{S}$ Services and ships $P^{\text{cha,max}}_{i}$ Permissible charger power at port $i$
$\mathcal{X}_{s}$ Feasible values of ship $s$ design variables $t^{\text{route}}_{s}$ Availability of service $s$
Decision variables $t^{\text{unit}}_{c}$ Loading duration of a unit of cargo $c$
$q_{ijcs}$ Quantity of cargo $c$ transported from port $i$ $U^{\text{min}}$ Required transportation service utility
to $j$ by ship $s$ during the planning horizon Functions
$N^{\text{trip}}_{s}$ Number of roundtrips completed in service $s$ $C^{\text{ship}}$ Ship acquisition and lifetime operation cost
$N^{\text{ship}}_{s}$ Number of ships deployed in service $s$ $E^{\text{leg}}_{ij}$ Ship electrical energy consumption
$P_{i}^{\text{cha}}$ Charger power at port $i$ sailing from port $i$ to $j$
$t^{\text{port}}_{ijs}$ Turnaround time the destination port $q^{\text{cap}}_{c}$ Ship capacity for cargo $c$
of leg $(i,j)$ of ship $s$ $V^{\text{GT}}$ Ship gross tonnage
$v_{ijs}$ Sailing speed in leg $(i,j)$ of ship $s$
$x^{\text{ship}}_{s}$ Vector of ship $s$ design variables
The model formulation is as follows:
Some text here that flows to the next line a bit moreminimize
$\sum_{s\in\mathcal{S}}\Biggl[N^{\text{ship}}_{s}C^{\text{ship}}(x^{\text{ship}}_{s})+\sum_{(i,j)\in\mathcal{L}_{s}}N^{\text{trip}}_{s}\Bigl(C_{j}^{\text{port}}V^{\text{GT}}(x^{\text{ship}}_{s})$ (1)
$+C_{j}^{\text{el}}E^{\text{leg}}_{ij}(x^{\text{ship}}_{s},v_{ijs})\Bigr)\Biggr]+\sum_{i^{\prime}\in\mathcal{N}}C^{\text{cha}}P_{i^{\prime}}^{\text{cha}},$
subject to
@@ 4.5 Model formulation
$\sum_{s:(i,j)\in\mathcal{A}_{s}}N^{\text{trip}}_{s}q_{ijcs}\leq q_{ijc}^{\text{dem}},\quad c\in\mathcal{C},\ (i,j)\in\mathcal{A},$ (8)
$\sum_{(i,j)\in\mathcal{L}_{s}}N^{\text{trip}}_{s}(t_{ijs}^{\text{transit}}+t_{ijs}^{\text{port}})\leq N^{\text{ship}}_{s}t_{s}^{\text{route}},\quad s\in\mathcal{S},$ (9)
$\max\{t_{ijs}^{\text{cargo}},t_{ijs}^{\text{cha}}\}\leq t_{ijs}^{\text{port}},\quad(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S},$ (10)
$\sum_{(i,j)\in\mathcal{A}}\sum_{s:(i,j)\in\mathcal{A}_{s}}\sum_{c\in\mathcal{C}}\alpha_{ijc}\log(N^{\text{trip}}_{s}q_{ijcs})\geq U^{\text{min}},$ (11)
$\quad x^{\text{ship}}_{s}\in\mathcal{X}_{s},\,s\in\mathcal{S},\quad P^{\text{cha}}_{i}\leq P^{\text{cha,max}}_{i},\,i\in\mathcal{N},$ (12)
where $t_{ijs}^{\text{transit}},\ldots,t_{ijs}^{\text{cha}}$ are to be substituted with the expressions
$t_{ijs}^{\text{transit}}=\frac{l_{ij}^{\text{leg}}}{v_{ijs}},$ (13)
$t_{ijs}^{\text{cargo}}=\max\{t_{ijsc}^{\text{load}}+t_{ijsc}^{\text{unload}}:c\in\mathcal{C}\},$ (14)
$t_{ijsc}^{\text{unload}}=t_{c}^{\text{unit}}\sum_{i^{\prime}\in\mathcal{N}^{-}_{js}\backslash\{j\}}q_{i^{\prime}jcs},$ (15)
$t_{ijsc}^{\text{load}}=\left\{\begin{array}[]{lr}t_{c}^{\text{unit}}\sum_{j^{\prime}\in\mathcal{N}_{s}\backslash\{j\}}q_{jj^{\prime}cs}&\text{if }j=d_{s}\\ t_{c}^{\text{unit}}\sum_{j^{\prime}\in\mathcal{N}^{+}_{js}\backslash\{j\}}q_{jj^{\prime}cs}&\text{otherwise,}\end{array}\right$. (18)
$t_{ijs}^{\text{cha}}=\frac{E^{\text{leg}}_{ij}(x^{\text{ship}}_{s},v_{ijs})}{P^{\text{cha}}_{j}}.$ (19)
The objective function (1) minimizes the total annualized cost. The four terms are the ship purchase and running cost, port charges, electricity cost, and shore charger installation cost. All these costs are scaled to one year.
The constraints are as follows:
  • 1.
    The capacity constraints (4) ensure that the ship capacity is not exceeded for any cargo type on the outbound legs. For the inbound legs (7), the capacity constraints take the same form, with the exception that the departure and arrival ports appear in a reverse order.
  • 2.
    The demand constraints (8) state that the demand of an origin-destination port pair is the upper bound on the total quantity of cargo transported by all the services that include both ports in their route.
  • 3.
    The upper bound on voyage duration is expressed in (9). The duration of a one-way voyage is the sum of times on sea (13) and port. Cargo unloading (15), loading (18), and charging (19) take place in parallel, and the operation that takes the longest determines the turnaround time (10).
  • 4.
    We require that the transportation network provide a minimum service level (11). Here, we use log-utility, which is a particular concave function with diminishing incremental value of cargo volume shipped in each origin-destination pair and cargo type.
4.6 Base ship design model
We now describe a particular simple ship sizing model for the ESFSMP defined in the previous section. This model uses a single sizing variable that scales the displacement of a reference ship. The accuracy is acceptable only for ship sizes within a small neighborhood of the reference ship.
The vector of design variables is $x^{\text{ship}}_{s}=(\nabla_{s},E^{\text{max}}_{s})$, where $\nabla$ is the displacement volume and $E^{\text{max}}_{s}$ is the battery capacity. Energy consumption on leg $(i,j)$ using speed $v_{ijs}$ is
$E_{ijs}^{\text{leg}}=t^{\text{transit}}_{ijs}P^{\text{ref}}\left(\frac{\nabla}{\nabla^{\text{ref}}}\right)^{2/3}\left(\frac{v_{ijs}}{v^{\text{ref}}}\right)^{3},$
where $P^{\text{ref}}$ is the propulsion power of a reference ship at $v^{\text{ref}}$ and displacement $\nabla^{\text{ref}}$. The feasible set is
$\mathcal{X}_{s}=\{x^{\text{ship}}_{s}:E_{ijs}^{\text{leg}}\leq E^{\text{max}}_{s},(i,j)\in\mathcal{L}_{s}\},\quad s\in\mathcal{S}.$ (20)
The functions in §4.4 take a monomial form, e.g., $q^{\text{cap}}_{c}(\nabla_{s})=\Psi_{c}^{(1)}(\nabla_{s})^{\Psi_{c}^{(2)}}$ with parameters $\Psi_{c}^{(1)},\Psi_{c}^{(2)}$.
4.7 Convex formulation
The ESFSMP described by (1) to (20) is a finite-dimensional, nonlinear, mixed-integer programming problem. It has, however, a special structure. The objective and the constraints are defined by log-convex functions. Hence, the problem is a mixed-integer log-convex optimization problem. As a result, it is transformed to a convex mixed-integer nonlinear programming problem using simple changes of variables. This transformation is exact, i.e., it does not involve any approximations. The transformed problem can be solved in a strong sense, with a guaranteed global optimality. For background on formulation techniques and solution methods for convex mixed-integer nonlinear programming, see Kronqvist et al. (2019).
In the following, we describe in detail the log-convex structure of the ESFSMP. Consider a particular log-convex function, called a posynomial, defined as
$f(x)=\sum_{k=1}^{K}c_{k}x_{1}^{a_{1k}}x_{2}^{a_{2k}}\cdots x_{n}^{a_{nk}},$ (21)
where $x\in\mathbb{R}^{n}_{++}$ is the $n$-vector of variables, $a_{1k},\ldots,a_{nk}\in\mathbb{R}$ and $c_{k}\in\mathbb{R}_{++}$ are constants and $K$ is the number of product terms. The ESFSMP can be expressed as a problem that minimizes a posynomial over a set defined by posynomial inequalities and the restriction that some number $P\leq n$ of variables are positive integers:
minimize $f_{0}(x)$ (22)
subject to $f_{m}(x)\leq 1,\quad m=1,\ldots,M,$
$x_{l}\in\mathbb{Z}_{\geq 1},\quad l=1,\ldots,P,$
where $f_{0},\ldots,f_{P}$ are posynomials. This problem becomes a convex mixed-integer nonlinear programming problem under a logarithmic change of variables, see A.
To see that the ESFSMP is defined by posynomials, consider the following reformulations. First, the capacity constraint (4), where $q^{\text{cap}}_{c}(\nabla_{s})=\Psi_{c}^{(1)}(\nabla_{s})^{\Psi_{c}^{(2)}}$, takes the posynomial inequality form (22) by a simple algebraic manipulation:
@@ 4.7 Convex formulation
The utility constraint (11) is equivalent to
$\prod_{(i,j)\in\mathcal{A}}\prod_{s:(i,j)\in\mathcal{A}_{s}}\prod_{c\in\mathcal{C}}(N^{\text{trip}}_{s}q_{ijcs})^{\alpha_{ijc}}\geq\exp{U^{\text{min}}}.$
The port turnaround time constraint (10) is equivalent to the inequalities
$t_{ijs}^{\text{port}}\geq t_{ijs}^{\text{load}},\quad(i,j)\in\mathcal{L}_{s},\,c\in\mathcal{C},\,s\in\mathcal{S},$
$t_{ijs}^{\text{port}}\geq t_{ijs}^{\text{unload}},\quad(i,j)\in\mathcal{L}_{s},\,c\in\mathcal{C},\,s\in\mathcal{S},$
$t_{ijs}^{\text{port}}\geq t_{ijs}^{\text{cha}},\quad(i,j)\in\mathcal{L}_{s},\,s\in\mathcal{S}.$
5 Physics-based ship design extension
Table 2: New decision variables for the extended ship design model. We use $s$ to index both services and ships because all the ships deployed on a service are identical.
Symbol Description
$L_{s}$ Hull length
$L^{\text{sup}}_{s}$ Superstructure length
$B_{s}$ Hull waterline breadth
$T_{s}$ Hull draft
$D_{s}$ Hull depth
$p_{s}^{\text{td}}$ Reference plate thickness
$h^{\text{batt}}_{s}$ Battery room height
$w^{\text{batt}}_{ks}$ Width of battery room $k$
$l^{\text{batt}}_{ks}$ Length of battery room $k$
$\tilde{h}^{\text{batt}}_{s}$ Battery room floor elevation from keel
$\tilde{l}^{\text{batt}}_{ks}$ Battery room $k$ front wall distance
from forward perpendicular
$\tilde{h}^{\text{roro}}_{s}$ First ro-ro deck elevation from keel
$E_{s}^{\text{batt}}$ Battery capacity
$t_{s}^{\text{cell}}$ Cell cycle life
$N_{s}^{\text{batt}}$ Number of battery replacements
$r_{ijs}^{\text{Cf}}$ Auxiliary variable for modeling frictional
resistance in sailing leg $(i,j)$
$r_{ijs}^{\text{Cr,std}}$ Auxiliary variable for modeling residual
resistance in sailing leg $(i,j)$
integration of hull wetted area
integration of hull steel area
This section describes a physics-based ship model that captures the key naval-architectural principles governing ship design. Although we focus on modeling a ro-pax ferry, the model applies to other ship types with minor modifications.
The model is expressed as a group of posynomial inequalities, meaning we can manipulate them to the form $f(x)\leq 1$, where $f$ is a posynomial function (21). Hence, the model we describe is compatible with the ESFSMP in §4.5. Table 2 lists the new decision variables that constitute the elements of the vector $x^{\text{ship}}$. The domain of all the variables is $\mathbb{R}_{++}$. All the constraints and the associated design variables hold for each $s\in\mathcal{S}$, but we drop the index for clarity. Note that some symbols used in §4.7 are redefined in this section.